the area of a circle is 200.96 cms. What is the diameter of the circle? please show work/explantion :D thanksz
-sandy

Answer by checkley71(8403)

AREA=piR^2 or 200.96=3.14r^2 or 200.96/3.14=r^2 or 64=r^2 or r=8